OpenStudy (anonymous):

Will Give Medal + Fan!!! Please solve for x. 5x-4/9=4 I believe it is 1.

OpenStudy (gorv):

5x-4/9= ????

OpenStudy (gorv):

i think u forgot that

OpenStudy (anonymous):

^^^

OpenStudy (anonymous):

I agree i'm really confused ?!? :/

OpenStudy (anonymous):

the problem is basically 5 times x minus 4 divided by 9

OpenStudy (anonymous):

can anyone help me?

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

The answer to the problem is 4

OpenStudy (anonymous):

I just need to figure out what x is...

OpenStudy (anonymous):

@AMYCARTER so its \[5x-4/9=4 ?\]

OpenStudy (anonymous):

yes @Fransisco_w

OpenStudy (anonymous):

What the equation is equal to is essential to answering the question, so include that next time.

OpenStudy (gorv):

it should be like 5 times x minus 4 divide by 9 equals to

OpenStudy (gorv):

otherwise u cannot solve it further

OpenStudy (anonymous):

The answer is 4 I added that just now

OpenStudy (gorv):

\[\frac{ 5x-4 }{ 9 }=4 ???\]

OpenStudy (anonymous):

@gorv I answered on my question :) sorry 4 taking so long multi tasking

OpenStudy (gorv):

plzz delete this comment ..dont spam others Q lol ..i m there